3Y 2 1. (20 points) Suppose that X and Y independent random variables. Let W 2x (a) Consider the following probability distri(b) Use your answers in part (a). If E(Y)=-3 and V(Y)= 1, what are E(W) and V (W)?
Answer #1

We have W=2x-34+2 os Mean of (x): {(x) = lixo) + (2.x0.7) + ( 3x0.3) 0-9 = = boy + 2.3 D3S ON van(x) = {(x) = ({(x))? or As,Page No. Date : As - sla.x + y) = ¿Caix+b) = a_{(x) + f(y)] anf(x) + b. I The variance of colis: vor (w) = vou (2x-3y + 2) =